OpenStudy (anonymous):

\[\begin{align*} 2x +3 & =4x +7 \\ -3 & = -3\\ \hline 2x & = 4x +4\\ -4x & = -4x \\ \hline -2x & =4 \\ x & = -2 \end{align*}\] At least the = signs line up. Still not exactly what I am looking for.

OpenStudy (anonymous):

\[\begin{align*} 2x \cancel{+3} & =4x +7 \\ \cancel{-3} & = -3\\ \hline 2x & = \cancel{4x} +4\\ -4x & = \cancel{-4x} \\ \hline -2x & =4 \\ x & = -2 \end{align*}\]

OpenStudy (anonymous):

@e.mccormick Is there a place holder I could use with the align* environment?

OpenStudy (e.mccormick):

What do you mean by a place holder?

OpenStudy (anonymous):

Something that will make an empty space the size of a character.

OpenStudy (e.mccormick):

In math mode, there are the assorted spaces, ~, \, \: \; \(|~|\,|\:| \;|\) where ~ is just short hand for \; width ads a space. There is also \quad and \qquad \( |\quad|\qquad |\). Any of those would put in dead space. Also, in an array or matrix, using & & skips a column. That should leave other things lined up... but I know I have had to toss a space in a few times to get things working the way I wanted.

OpenStudy (anonymous):

\[\begin{align*} 2x \cancel{+3} & =~4x +7 \\ \cancel{-3} & = ~~~~~-3\\ \hline 2x & = \cancel{4x} +4\\ -4x & = \cancel{-4x} \\ \hline -2x & =4 \\ x & = -2 \end{align*}\] That will work. Thanks.

OpenStudy (ybarrap):

$$\large \text{"\hspace{x pt}" Seemed to do work well for this problem:}$$ \[\begin{align*} 2x \cancel{+3} & =~4x +7 \\ \cancel{-3} & =\hspace{18 pt}-3\\ \hline 2x & = \cancel{4x} +4\\ -4x & = \hspace{-7 pt} \cancel{-4x} \\ \hline -2x & =\hspace{29 pt}4 \\ x & = \hspace{21 pt} -2 \end{align*}\\ \\ \Large \text{Here is the }\LaTeX \Large \text{ I used:} \] $$ \begin{align*} 2x \cancel{+3} & =~4x +7 \\ \cancel{-3} & =\hspace{18 pt}-3\\ \hline 2x & = \cancel{4x} +4\\ -4x & = \hspace{-7 pt} \cancel{-4x} \\ \hline -2x & =\hspace{29 pt}4 \\ x & = \hspace{21 pt} -2 \end{align*} }$$

OpenStudy (e.mccormick):

@ybarrap No need to post the \(\LaTeX\) you used. We can right click it. \(\ddot \smile\) If the goal was to keep them in columns, then the easy way is to just use those columns. \(\begin{array}{rrcrr} 2x &\cancel{+3} & = &4x &+7 \\ & \cancel{-3} & =& &-3\\ \hline 2x & & =& \cancel{4x} &+4\\ -4x && =& \cancel{-4x} \\ \hline -2x && =& &4 \\ x & & = & &-2 \end{array}\) But personally I find the spread a bit much. Which is where the column width setting is nice, but does not work here.... Collapsing spaces is OK, but you are bound to miss something... \(\begin{array}{rrcrr} \! 2x \! &\cancel{+3} & \! = \! &4x &\! +7 \! \\ \! & \! \cancel{-3} &\! =\! &\! &\! -3\! \\ \hline \! 2x &\! &\! =\! & \cancel{4x}\! &\! +4\! \\ \! -4x &\! & \! =\! &\! \cancel{-4x}\! \\ \hline \! -2x &\! &\! =\! & \! \! &\! 4\! \\ \! x\! &\! &\! = \! &\! \! &\! -2\! \end{array}\) And you will never quite get the spacing right because the system was designed to tweak spacing independent of what you enter. So as you adjust spacing, so does it, and they rarely match perfectly. Another option is to just break it into pieces and let it shrink as you go, not worrying so much about it. \(\begin{align*} 2x \cancel{+3} & =~4x +7 \\ \cancel{-3} & = ~~~~~-3\\ \hline \end{align*}\) \(\begin{align*} 2x & = \cancel{4x} +4\\ -4x & = \cancel{-4x} \\ \hline \end{align*}\) \(\begin{align*} -2x & =4 \\ x & = -2 \end{align*}\) Or, use the implies arrow. \[2x +3 =4x +7 \implies\\ 2x \cancel{+3-3} =4x +7-3\implies \\ 2x = 4x +4\implies\\ 2x-4x = \cancel{4x-4x} +4\implies\\ -2x =4 \implies \\ \cancel{\tfrac{-2}{-2}} x = \tfrac{4}{-2}\implies \\ x=-2\] And if you want to space that to make it clearer, don't do it in one block. \(2x +3 =4x +7 \implies\) \(2x \cancel{+3-3} =4x +7-3\implies \) \(2x = 4x +4\implies\) \(2x-4x = \cancel{4x-4x} +4\implies\) \(-2x =4 \implies \) \(\cancel{\tfrac{-2}{-2}} x = \tfrac{4}{-2}\implies \) \(x=-2\)

OpenStudy (ybarrap):

nice suggestions, thanks!

OpenStudy (anonymous):

Great ideas! Thanks!

OpenStudy (e.mccormick):

Just learned about this page: http://www.onemathematicalcat.org/MathJaxDocumentation/TeXSyntax.htm Lists all the MathJax stuff... but is a little off. Says cancel won't work, which we know it does. However, it does talk a bit about alignat, which is supposed to be more compressed than align.

OpenStudy (e.mccormick):

aligned \[\begin{aligned} 2x\cancel{+3} & = 4x +7 \\ \cancel{-3} & = -3\\ \hline 2x & = \cancel{4x} +4\\ -4x& = \cancel{-4x} \\ \hline -2x & =4 \\ x & = -2 \end{aligned}\] hmm.... Lots of options on that site.

OpenStudy (e.mccormick):

\(\huge\rlap{\,\;\text{This question rates as:}}\bbox[5pt,#AA00AA ,border:3px solid #2200EE ]{\color{#FFFF00 }{\text{This question rates as:}}}\) \(\Huge \bbox[2pt,red,border:3px solid gold]{\color{gold}{★ }}\bbox[2pt,red,border:3px solid gold]{\color{gold}{★ }}\bbox[2pt,red,border:3px solid gold]{\color{gold}{★ }}\bbox[2pt,red,border:3px solid gold]{\color{gold}{★ }}\) On my personal scale... Hehe. =)

OpenStudy (e.mccormick):

Thought you might like those. Just for fun.

OpenStudy (ybarrap):

$$\huge\bbox[5pt,#AA00AA ,border:3px solid #2200EE ]{\color{#FFFF00 }{\frak{Pretty ~Cool!!}}}$$

OpenStudy (e.mccormick):

\(\huge\rlap{\,\;\frak{Pretty~Cool!!}}\bbox[5pt,#AA00AA ,border:3px solid #2200EE ]{\color{#FFFF00 }{\frak{Pretty~Cool!!}}}\;\normalsize\leftarrow\) With shadow. Gets a nice embossed effect because of the white to the right.